Suppose that Lil John Industries equity is currently selling for $46 per share and that 3.9 million shares are outstanding. The firm also has 69,000 bonds outstanding, which are selling at 102 percent of par. Assume Lil John was considering an active change to its capital structure so that the firm would have a (D/E) of 1.5. Which type of security(stocks or bonds) would it need to sell to accomplish this? sell bonds and buy back stock sell stocks and buy back bonds How much would the firm have to sell? (Enter your answer in dollars not in millions. Do not round intermediate calculations and round your final answer to 2 decimal places.) Selling amount
Answer #1

Equity Value = 46 * 3,900,000 = 179,400,000
Debt Value = 102* 69,000 = 7,038,000

1 Option a sell stocks and buy back bonds

2. Debt/Equity = 1.5 = (7038000+x)/(179400000-x)
=  1.5*179400000 -1.5x = 7038000+x
2.5x = 1.5*179400000 - 7038000
x = 104,824,800

The firm should sell = 104.824,800

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well